a series rl is connected to an ac source. voltage across the indulgence is 30 v voltage across the resistor is 40 v. what is the effective voltage of the source
A.70V b. 50V c. 10 V d. 30V
1
Expert's answer
2018-05-08T10:58:08-0400
If voltage across inductance and resistor (30 V and 40 V) is effective voltage:
U_source^effective=30+40=70 V
If voltage across inductance and resistor (30 V and 40 V) is amplitude voltage:
U_source^effective=30/√2+40/√2=50 V